OpenStudy (anonymous):

R=log(I/Io) R is the Richter number of the earthquake, I is the intensity of the earthquake, and Io is the intensity of the minimum detectable earthquake. Please help! An earthquake has a Richter number of 4.2. How many times more intense than the minimum detectable earthquake is the earthquake? Round your answer to the nearest whole number.

OpenStudy (anonymous):

\[R = 4.2\] and the answer you are looking for is: \[\frac{I}{I_0}\] You want a ratio of the intensity to the minimum detectable intensity (you want the value of that fraction). so you want to solve: \[4.2 = \log\bigg(\frac{I}{I_0}\bigg)\] for the fraction. your answer should be something like \[\frac{I}{I_0} = number\]